Average varible caost $10 at a quantity of 100, and minmum average cost $15 at a quantity of 200 hamburgers. Fixed cost is $300. What would the average Variable cost be for a quantity of 200?
1
Expert's answer
2012-11-14T08:33:13-0500
AFC = FC/Q
AC = AVC + AFC = AVC + FC/Q,
AVC = AC - FC/Q = 15 - 300/200 = $13.5